OpenStudy (anonymous):

What are the retricemptotes of the hyperbola: (x-1)^(2)/(25)-(y-4)^(2)/(16)=1

OpenStudy (amistre64):

bx/a where b and a are the sqruare roots of your denoms

OpenStudy (amistre64):

since the x spot is bigger; that isthe width of the rectangular asymptotes of the hyperbola

OpenStudy (amistre64):

-5,5 the height of this rectangular thing is -4,4

OpenStudy (amistre64):

i just noticed its off centered lol

OpenStudy (amistre64):

(x-1)^2 (y-4)^2 -------- - ------- = 1 a^2=25 b^2=16 if this thing were centered wed have the rectangle asymptote at: y = (4/5) x but lets put it back now

OpenStudy (amistre64):

(y-4) = (4/5)(x-1) y = (4/5)x - 4/5+20/5 y = (4/5)x + 16/5

OpenStudy (amistre64):

and the other is just the negative slope of that i believe

OpenStudy (anonymous):

so is (4/5)x+16/5 the retricepmtote?

OpenStudy (amistre64):

if i did it right, it should be one of them; th eother would be: y = -(4/5)x + 16/5

OpenStudy (amistre64):

y = (b/a)x is the normal asymptote; but since this is off centered, we gotta adjust for that discrepency

OpenStudy (amistre64):

gotta get to calc class and learn limits lol.... have fun ;)

OpenStudy (anonymous):

it can have multiple retricepmtotes, so would it be wrriten as (4/5)x, and 16/5?